You answered one question for me, and I have another just so I get the hang of how to set it up. if 8 oz of chili has 17 grams of protein how many protein in 1 oz?
Thank you
Shelley

Answer by stanbon(75887)

if 8 oz of chili has 17 grams of protein how many protein in 1 oz
-------------
Proportion:
protein/weight = protein/weight
----------------------------
x/1oz = 17/8oz
------
x = 1(17/8)
----
x = 2 1/8 grams of protein
===============================
Cheers,
Stan H.
============